A coffee distributor needs to mix a(n) Ethiopian coffee blend that normally sells for $8.10 per pound with a Breakfast coffee blend that normally sells for $14.30 per pound to create 80 pounds of a coffee that can sell for $12.67 per pound. How many pounds of each kind of coffee should they mix?
A) Write an equation using the information as it is given above that can be solved to answer the question. Use
x as your variable to represent the quantity of Ethiopian coffee blend.
Equation:
B) Answer: They must mix
?pounds of the Ethiopian Blend
?pounds of the Breakfast Blend.
Round your answers to the nearest whole number of pounds.

Answer by mananth(16946)


Ethiopian coffee blend that normally sells for $8.10 per pound . let it be x pounds
Breakfast coffee blend that normally sells for $14.30 per pound let it be y pounds
to create 80 pounds of a coffee(x+y) that can sell for $12.67 per pound.
Quantity
x+y=80-----------------------(1)
8.10x+14.30y = 12.67(x+y)
8.10x + 14.30y = 12.67x+12.67y
rearrange.
4.57x-1.63y=0--------------------(2)
1.00 x + 1.00 y = 80.00


4.57 x + -1.630 y = 0.00 .............2
Eliminate y
multiply (1)by 1.63
Multiply (2) by 1.00
1.63 x 1.63 y = 130.40
4.57 x -1.63 y = 0.00
Add the two equations
6.20 x = 130.40
/ 6.20
x = 21.03
plug value of x in (1)
1.00 x + 1.00 y = 80.00
21.03 + 1.00 y = 80.00
1.00 y = 58.97
y = 58.97
Ans x = 21.03
y = 58.97
21.03 pounds Ethopian coffee
58.97 pounds Breakfast coffee